Given that a-3b/a+2b=3, find the value of a+3b/a-2b. Express your answer as an integer or common or improper fraction reduced to lowest terms. Do not use decimals.

Answer by Boreal(15235)

a-3b/a+2b=3,
a-3b=3a+6b
-2a=9b
a= (-9/2)b
============
(a+3b)/(a-2b)
(-9b/2) + (6b/2)/(-9b/2)-(4b/2)
(-3b/2)/(-13b/2)
3/13